# Residual Symmetries and BRST Cohomology of Schwarzschild in the Kerr-Schild Double Copy

### v5

@@ 3.5.1 Nilpotency on Fields
Swapping $b\leftrightarrow c$ in the structure constants ${f_{ba}}^{c}$ gives:
$\mathcal{Q}^{2}\Psi=\frac{1}{2}{f_{ab}}^{c}c^{a}c^{b}\mathcal{L}_{K_{c}}\Psi+\frac{1}{2}{f_{ab}}^{c}c^{b}c^{a}\mathcal{L}_{K_{c}}\Psi=\frac{1}{2}{f_{ab}}^{c}(c^{a}c^{b}+c^{b}c^{a}).$ (102)
@@ 3.5.2 Nilpotency on Ghosts
$\mathcal{Q}^{2}c^{a}=-\frac{1}{2}{f_{bc}}^{a}\mathcal{Q}(c^{b})c^{c}+\frac{1}{2}{f_{bc}}^{a}c^{b}\mathcal{Q}(c^{c})=\frac{1}{4}{f_{bc}}^{a}{f_{de}}^{b}c^{d}c^{e}c^{c}-\frac{1}{4}{f_{bc}}^{a}{f_{de}}^{c}c^{b}c^{d}c^{e}.$ (103)
Anticommuting $c^{c},c^{e}$, then $c^{c},c^{d}$, and relabeling $b\leftrightarrow c$ in the first term yields:
$\mathcal{Q}^{2}c^{a}=-\frac{1}{4}{f_{cb}}^{a}{f_{de}}^{c}c^{b}c^{d}c^{e}-\frac{1}{4}{f_{bc}}^{a}{f_{de}}^{c}c^{b}c^{d}c^{e}.$ (104)
Finally, permuting $b\leftrightarrow c$ in ${f_{cb}}^{a}$ introduces a minus sign to the first term and the terms cancel. Hence, $\mathcal{Q}^{2}c^{a}=0$. Therefore, $\mathcal{Q}$ is nilpotent. $\square$
@@ 3.5.3 Action on the Kerr-Schild Metric
In the Schwarzschild Kerr-Schild background, the metric takes the form of (52), and the residual generators $K_{a}\in{K_{0},K_{i}}$ are precisely the Killing vectors of the spacetime. Consequently, the BRST charge acts trivially on the metric:
@@ 4 Conclusion and Discussion
Acknowledgments
I would like to thank Dr. Silvia Nagy for her guidance and insightful feedback throughout this project. I am also grateful to my previous thesis advisor, Dr. Xavier Calmet, who was gracious enough to endorse this paper. Moreover, I am appreciative of my colleagues and collaborators, including Dr. Martin Bauer and Dr. Edwin Hach, for valuable discussions and support.

### v8

@@ 3.5.1 Nilpotency on Fields
Permuting $a\leftrightarrow b$ in the structure constants ${f_{ba}}^{c}$ gives:
$\mathcal{Q}^{2}\Psi=\frac{1}{2}{f_{ab}}^{c}c^{a}c^{b}\mathcal{L}_{K_{c}}\Psi+\frac{1}{2}{f_{ab}}^{c}c^{b}c^{a}\mathcal{L}_{K_{c}}\Psi=\frac{1}{2}{f_{ab}}^{c}(c^{a}c^{b}+c^{b}c^{a})\mathcal{L}_{K_{c}}\Psi.$ (102)
@@ 3.5.2 Nilpotency on Ghosts
$\mathcal{Q}^{2}c^{a}=-\frac{1}{2}{f_{bc}}^{a}\mathcal{Q}(c^{b})c^{c}+\frac{1}{2}{f_{bc}}^{a}c^{b}\mathcal{Q}(c^{c})=\frac{1}{4}{f_{bc}}^{a}{f_{de}}^{b}c^{d}c^{e}c^{c}-\frac{1}{4}{f_{bc}}^{a}{f_{de}}^{c}c^{b}c^{d}c^{e},$ (103)
by the graded Leibniz rule. Anticommuting $c^{c},c^{e}$, then $c^{c},c^{d}$, and relabeling $b\leftrightarrow c$ in the first term yields:
$\mathcal{Q}^{2}c^{a}=\frac{1}{4}{f_{cb}}^{a}{f_{de}}^{c}c^{b}c^{d}c^{e}-\frac{1}{4}{f_{bc}}^{a}{f_{de}}^{c}c^{b}c^{d}c^{e}.$ (104)
By the Jacobi identity
${f_{bc}}^{a}{f_{de}}^{c}+{f_{cd}}^{a}{f_{eb}}^{c}+{f_{ce}}^{a}{f_{bd}}^{c}=0\implies{f_{bc}}^{a}{f_{de}}^{c}=-[{f_{cd}}^{a}{f_{eb}}^{c}+{f_{ce}}^{a}{f_{bd}}^{c}],$ (105)
we are free to write (104) as
$\mathcal{Q}^{2}c^{a}=\frac{1}{4}\begin{pmatrix}{f_{cb}}^{a}{f_{de}}^{c}+{f_{cd}}^{a}{f_{eb}}^{c}+{f_{ce}}^{a}{f_{bd}}^{c}\end{pmatrix}c^{b}c^{d}c^{e}.$ (106)
The sum of structure constants satisfies the Jacobi identity as well, so (106) vanishes. Hence, $\mathcal{Q}^{2}c^{a}=0$. Therefore, $\mathcal{Q}$ is nilpotent. $\square$
@@ 3.5.3 Action on the Kerr-Schild Metric
In the Schwarzschild Kerr-Schild background, the metric takes the form of (52), and the residual generators $K_{a}\in\{K_{0},K_{i}\}$ are precisely the Killing vectors of the spacetime. Consequently, the BRST charge acts trivially on the metric:
@@ 4 Conclusion and Discussion
Acknowledgements
I would like to thank Dr. Silvia Nagy for providing invaluable insight and expertise on the double copy literature. I couldn’t have written this paper without her input, corrections, and constructive criticism.